OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

@!MilitaryLad!

OpenStudy (studygurl14):

Equation of a line: y = mx + b m = slope So, first step, what can you plug in for m?

OpenStudy (anonymous):

-3

OpenStudy (studygurl14):

Also, remember that coordinates are (x,y), so x = 2, and y = -5

OpenStudy (studygurl14):

Correct @<3bae<3

OpenStudy (amorfide):

If you keep in mind the formula for finding the equation of a line y-y1=m(x-x1) where y1 is the y ordinate, and x1 is the x ordinate, and m is the gradient (slope) substitute in your values and simplify

OpenStudy (anonymous):

i cant do this one.....:(